The following data represent the results of inspection of bearings supplied by three different manufacturers Company A Company B Company C Satisfactory Oversized Undersized 3000 880 200 2500 100 150 1500 120 250 Find the probability that a) a defective part is oversized (b) Company A supplied a defective part, and (c) a part is both defective and from Company C
Answer #1

from above:

A B C Total
Satisfactory 3000 2500 1500 7000
Oversized 880 100 120 1100
Undersized 200 150 250 600
Total 4080 2750 1870 8700

a)

P(defective part is oversized)=1100/(1100+600)=11/17 (As 1100 oversized part from 1700 defective)

b)

P(defective part given from Company A)=(880+200)/4080=1080/4080=9/34

c)

P(defective and from company C)=(120+250)/8700=370/8700=37/870